Given the following ANOVA table:

Source DF SS MS F
Regression 1 1,190.0 1,190.00 38.90
Error 17 520.0 30.59
Total 18 1,950.0
  1. Determine the coefficient of determination. (Round your answer to 3 decimal places.)

  1. Assuming a direct relationship between the variables, what is the correlation coefficient? (Round your answer to 2 decimal places.)

  1. Determine the standard error of estimate. (Round your answer to 2 decimal places.)
Answer #1

a. The co-efficient of determination, = 1 - (SSE/SST) = 1 - (520/1950) = 1 - 0.2667 = 0.7333 (Ans).

b. The correlation co-efficient, r = = 0.8563 (Ans).

c. The standard error of the estimate = = = 5.5308 (Ans).